Grade 12th passMechanics

The electric potential in a region of space is given by V = (5x – 7x 2 y + 8y 2 + 16yz – 4z) volt. The y-component of the electric field at the point (2, 4, -3) is A) 7 volt/ m B) 12 volt/ m C) 16 volt/ m D) 31 volt/ m

aswin a
7 Years agoGrade 12th pass

2 Answers

Arun
7 Years ago
Take partial derivatives with respect to each coordinate: 
V = 5x -3x^2y + 2yz^2 
Ex = -∂V/∂x = -[5 - 6xy] = -5 + 6xy 
Ey = -∂V/∂y = -[-3x^2 + 2z^2] = 3x^2 - 2z^2 
Ez = -∂V/∂z = -4yz
abc ytb
4 Years ago
Ey = - [ 0 - 7x2 + 16y + 16z ]
     = - [ - 7 * 4 + 16 * 4 + 16 * (-3) ]
     =  28 - 64 + 48
     = 12 v/m
...........................................................................................................